OpenStudy (anonymous):

Assume that x and y are both differentiable functions of t and find the required values of dy/dt and dx/dt. y = √x (a) Find dy/dt, given x = 16 and dx/dt = 7. dy/dt = (b) Find dx/dt, given x = 64 and dy/dt = 8. dx/dt =

OpenStudy (anonymous):

y = √x (a) Find dy/dt, given x = 16 and dx/dt = 7. dy/dt = (b) Find dx/dt, given x = 64 and dy/dt = 8. dx/dt =

OpenStudy (anonymous):

\[\frac{dy}{dt}=\frac{1}{2\sqrt{x}}\frac{dx}{dt}\]

OpenStudy (anonymous):

hey satelitte:)

OpenStudy (anonymous):

put in x = 16 and dx/dt=7 to get your answer!

OpenStudy (anonymous):

you get \[\frac{dy}{dt}=\frac{1}{2\sqrt{16}}\times 7\]

OpenStudy (anonymous):

or \[\frac{dy}{dt}=\frac{7}{8}\]

OpenStudy (anonymous):

for the second one put \[8=\frac{1}{2\sqrt{64}}\times \frac{dx}{dt}\] and solve

OpenStudy (anonymous):

i get \[\frac{dx}{dt}=128\]

OpenStudy (anonymous):

yes!!!:)..ur a complete life saver today you have no idea.god bless you!:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

just pass your class ok?

OpenStudy (anonymous):

yeh im trying..ur helping me out and i appreciate it alot

OpenStudy (anonymous):

i absolutely loved jennifer's body. laughed my @ss off. maybe favorite horror movie of 2009